OpenStudy (anonymous):

Find the area of a rectangle if the perimeter is equal to 38 in and the length is 10.

OpenStudy (raden):

u must find its width, first

OpenStudy (anonymous):

how do u do tht

OpenStudy (raden):

do u know formula of the perimeter a rectangle

OpenStudy (anonymous):

no can you please tell me

OpenStudy (raden):

perimeter = 2(l+w) with l=length, w=width

OpenStudy (raden):

so, u can make an equation : 2(10+w) = 38 solve for w

OpenStudy (anonymous):

so it will be 2(9.5+10)= 39

OpenStudy (raden):

nopes.. it should be 2(10+w) = 38 now, solve for w

OpenStudy (anonymous):

w=18

OpenStudy (raden):

nopes.. 2(10+w) = 38 both sides divided by 2, what u get ...

OpenStudy (anonymous):

2 * 10 + w = 38 20+w=38 w=18

OpenStudy (raden):

no, if u want multiply in bracket by 2, u must use distributive property : a(b+c) = a*b + a*c so, 2(10+w) = 2*10 + 2w = 20 + 2w

OpenStudy (raden):

now, u have 20+2w=38

OpenStudy (anonymous):

but that is not one of the choices I have on my homework

OpenStudy (raden):

yeah, because this is not clear. after u found the velue of w, next u can determine the area by using the formula A=l*w

OpenStudy (anonymous):

80?

OpenStudy (anonymous):

90*

OpenStudy (raden):

well, look this : 20 + 2w = 38 move 20 to right side, becomes 2w = 38-20 2w = 18 w = 18/2 = 9 so, the area (A) = l * w = 10 * 9 = ...

OpenStudy (anonymous):

yup its 90 thank u so much

OpenStudy (raden):

yup, very welcome